OpenStudy (anonymous):

Company decides to take random sample of 10 cell phones from production run of 2500 cell phones. Mean number of faulty cell phones is 1.6/day, what is probability of faulty cell phones produced tomorrow (one day has no bearing on the next). How to solve?

OpenStudy (anonymous):

Correction: What is the probability that at least 1 faulty cell phone will be produced tomorrow*

OpenStudy (anonymous):

I'm trying to solve this, but everything I'm trying comes out with 400 and i don't think that would be correct.

OpenStudy (anonymous):

Similar example: mean = 4.8, prob at least 3 faulty cell phones is 0.8575

OpenStudy (anonymous):

Hopefully this will help. I never liked this subject. Haha http://oscience.info/math-formulas/probability-formulas/

OpenStudy (anonymous):

It's a bit more advanced

OpenStudy (anonymous):

Yes, what math are you taking? You're in college, right?

OpenStudy (anonymous):

Yes - Business Statistics

OpenStudy (anonymous):

I solved it: 1 - [p(x=0) + p(x=1) + p(x=2)] Formula for each p is (e^(-lambda) x lambda)/x!